1. The 1933 earthquake in Japan registered 8.9 on the Richter scale. a). What would be the Richter scale reading of an earthquake that is only 0.01 times as intense as the earthquake in Japan? b). The 1985 earthquake in Peru registered 7.7 on the Richter scale. How many times more intense was the earthquake in Japan compared to the Peruvian quake? Round your answer to one decimal place.
Answer by LinnW(1048)

a) 0.01 times as intense, 10^(-2) * 10(8.9) = 10(6.9)
so 6.9 would be 0.01 times as intense.
b) 8.9 - 7.7 so 10^(-2.2) times = 0.00631 = 0.01 to one position